OpenStudy (animelover91804):

Find the mean of this set of numbers. 9, –6, –8, 11, –4, 10 A. 12 B. 48 C. 2 D. 8

OpenStudy (smilie2015):

Count the number N of data points in your data set. Sum up the total of your data set; add them all together. Average = Mean = Sum/N.

OpenStudy (smilie2015):

=9+( –6)+ (–8) +11+(–4)+10 =12 =12/6 =2